A military supply plain is on point to a mission to drop supplies in. Its flying 1200 m above the ground at 723 m/s to avoid enemy fire. To drop its pay load (800 kg) on the correct target, given the plains speed and altitudeWhat will the final angle just before it hits the ground?
Answer #1

Solution > 73x4s 1200 m for the projectile at time it y = y; - 12gt² at t y=0 0 = 1200 -29 2x1200 = 15.668. 9-81 Vy = -97 =

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions.